# Wavefront Curvature and Transverse Atomic Motion in Time-Resolved Atom Interferometry: Impact and Mitigation

### v1

@@ IV.2 Learning the phase perturbation from position-resolved data
An example of a position-resolved measurement protocol is phase-shear readout [53], in which a linear phase gradient is imprinted onto the atom cloud using a tilted mirror, creating spatial interference fringes whose offset encodes the interferometric phase. This allows for single-shot phase estimation and makes phase-shear readout especially well suited for time-resolved atom interferometry [25]. In what follows, we focus on this scheme because it requires only one simulation per phase-shift estimate, but our conclusions could in principle apply to other measurement protocols that resolve final atomic positions.
@@
where expressions for the coefficients $\beta_{i}$ are given in Table 2 and derived in Appendix C, for the case $\bm{\theta}=(\mu_{x},\mu_{v_{x}})^{T}$ (expressions in terms of initial COM coordinates are readily obtained by substituting $\mu_{x}=\mu_{x_{0}}+\mu_{v_{x0}}t_{\text{det}}$ and $\mu_{v_{x}}=\mu_{v_{x0}}$). However, if the cloud is localized around $\mu_{x}$, the position-resolved phase will appear approximately linear in $x$, such that, assuming that the cloud is Gaussian-distributed in the transverse directions at detection time, the distribution can be accurately fit to the model [53]
@@
Interestingly, this also implies that, if variations in $\mu_{x_{0}}$ and $\mu_{v_{x0}}$ are experimentally controlled and $\Delta\varphi_{0}$ remains constant shot-to-shot, observed variations in $\Delta\varphi$ and $\kappa$ can be used as regression parameters (together with an appropriate model, such as the simulations presented here) to estimate the underlying beam parameters. This approach could enable in-situ beam characterisation, providing a pathway toward systematic error estimation in precision measurements, for instance in determinations of the fine-structure constant [6, 7, 8].
@@ IV.3.2 Robustness to other couplings
The present analysis assumes phase-gradient fluctuations arise solely from coupling between transverse motion and curved wavefronts. In practice, other mechanisms may contribute. For example, the Coriolis force can imprint a phase gradient across the atom cloud [47]. While future long-baseline interferometers are expected to employ Coriolis compensation schemes, residual compensation errors could still introduce phase-gradient noise.
@@ V Conclusion
We applied our results to estimate phase noise in 100-m and 1-km gradiometers modelled after upcoming long-baseline atom interferometry experiments such as AION [24] and MAGIS-100 [25]. We found that to achieve sub-$10^{-5}$ rad phase noise levels (the target sensitivity for these experiments) for a 100-m detector with $w_{0}=1$ cm, stability at the micron and sub-micron-per-second levels is necessary in the HEHN configuration, while tens-of-microns and few-microns-per-second stability suffice for the LELN case. This imposes stringent requirements on source preparation stability in future long-baseline atom interferometry experiments, and could necessitate significant improvements over current performances.
@@ Appendix B Monte Carlo simulation
To validate the analytical models derived in this work, we performed Monte Carlo (MC) simulations of a large momentum transfer Mach-Zehnder atom interferometer. Each simulation evaluates the total interferometric phase shift $\Delta\varphi$, given by equation (1), for a large ensemble of individual trajectories with initial positions and velocities drawn from a Gaussian distribution, and evolved by numerically solving Hamilton’s equations of motion. The propagation phase in equation (2) is computed via numerical integration by evaluating the classical action along each trajectory. The quantum evolution of the internal states is modelled by solving the time-dependent Schrödinger equation for a Gaussian wave packet with two internal degrees of freedom in a plane wave with phase offset $\phi(\bm{x}_{i})$, where $\bm{x}_{i}$ is the wave packet’s central coordinate at the interaction time, following the approach in Refs. [54, 38]. The simulations use a position-dependent Rabi frequency to represent arbitrary laser intensity profiles. At the end of the interferometry sequence, trajectories that overlap in phase space and share the same internal state interfere, forming an effective interferometry port. The probability of detecting an atom at a given port is:
@@
In the one-dimensional position-resolved case, all phase-space coordinates but $x$ are integrated over. For the sake of conciseness, we denote the integrated coordinates $\bm{\xi}_{I}={(y,z,v_{x},v_{y},v_{z})}^{T}$. Using properties of normal distributions (see Ref. [55]) it can then be shown that the integrated coordinates conditioned on $x$ are also normally distributed, i.e., $\bm{\xi}_{I}|x\sim\mathcal{N}(\bm{\mu}_{\xi_{I}|x},\bm{\Sigma}_{\xi_{I}|x})$, with

### v2

@@ IV.2 Learning the phase perturbation from position-resolved data
An example of a position-resolved measurement protocol is phase-shear readout [Sugarbaker_2013], in which a linear phase gradient is imprinted onto the atom cloud using a tilted mirror, creating spatial interference fringes whose offset encodes the interferometric phase. This allows for single-shot phase estimation and makes phase-shear readout especially well suited for time-resolved atom interferometry [Abe_2021]. In what follows, we focus on this scheme because it requires only one simulation per phase-shift estimate, but our conclusions could in principle apply to other measurement protocols that resolve final atomic positions.
@@
where expressions for the coefficients $\beta_{i}$ are given in Table 2 and derived in Appendix C, for the case $\bm{\theta}=(\mu_{x},\mu_{v_{x}})^{T}$ (expressions in terms of initial COM coordinates are readily obtained by substituting $\mu_{x}=\mu_{x_{0}}+\mu_{v_{x0}}t_{\text{det}}$ and $\mu_{v_{x}}=\mu_{v_{x0}}$). However, if the cloud is localized around $\mu_{x}$, the position-resolved phase will appear approximately linear in $x$, such that, assuming that the cloud is Gaussian-distributed in the transverse directions at detection time, the distribution can be accurately fit to the model [Sugarbaker_2013]
@@
Interestingly, this also implies that, if variations in $\mu_{x_{0}}$ and $\mu_{v_{x0}}$ are experimentally controlled and $\Delta\varphi_{0}$ remains constant shot-to-shot, observed variations in $\Delta\varphi$ and $\kappa$ can be used as regression parameters (together with an appropriate model, such as the simulations presented here) to estimate the underlying beam parameters. Complementing existing in-situ characterisation techniques, such as probing the spatial distribution of wavevectors using Bose-Einstein condensates [gaudout2025probingspatialdistributionkvectors], this approach would provide an alternative pathway toward systematic error estimation in precision measurements, for instance in determinations of the fine-structure constant [Bouchendira2010, Parker, Morel2020].
@@ IV.3.2 Robustness to other couplings
The present analysis assumes phase-gradient fluctuations arise solely from coupling between transverse motion and curved wavefronts. In practice, other mechanisms may contribute. For example, the Coriolis force can imprint a phase gradient across the atom cloud [PhysRevLett.111.083001]. While future long-baseline interferometers are expected to employ Coriolis compensation schemes, residual compensation errors could still introduce phase-gradient noise.
@@ V Conclusion
We applied our results to estimate phase noise in 100-m and 1-km gradiometers modelled after upcoming long-baseline atom interferometry experiments such as AION [Badurina_2020] and MAGIS-100 [Abe_2021]. We found that to achieve sub-$10^{-5}$ rad phase noise levels (the target sensitivity for these experiments) for a 100-m detector with $w_{0}=1$ cm, stability at the micron and sub-micron-per-second levels is necessary in the HEHN configuration, while tens-of-microns and few-microns-per-second stability suffice for the LELN case. This imposes stringent requirements on source preparation stability in future long-baseline atom interferometry experiments, and could necessitate significant improvements over current performances.
@@ Appendix B Monte Carlo simulation
To validate the analytical models derived in this work, we performed Monte Carlo (MC) simulations of a large momentum transfer Mach-Zehnder atom interferometer. Each simulation evaluates the total interferometric phase shift $\Delta\varphi$, given by equation (1), for a large ensemble of individual trajectories with initial positions and velocities drawn from a Gaussian distribution, and evolved by numerically solving Hamilton’s equations of motion. The propagation phase in equation (2) is computed via numerical integration by evaluating the classical action along each trajectory. The quantum evolution of the internal states is modelled by solving the time-dependent Schrödinger equation for a Gaussian wave packet with two internal degrees of freedom in a plane wave with phase offset $\phi(\bm{x}_{i})$, where $\bm{x}_{i}$ is the wave packet’s central coordinate at the interaction time, following the approach in Refs. [antoine2004, antoine2006]. The simulations use a position-dependent Rabi frequency to represent arbitrary laser intensity profiles. At the end of the interferometry sequence, trajectories that overlap in phase space and share the same internal state interfere, forming an effective interferometry port. The probability of detecting an atom at a given port is:
@@
In the one-dimensional position-resolved case, all phase-space coordinates but $x$ are integrated over. For the sake of conciseness, we denote the integrated coordinates $\bm{\xi}_{I}={(y,z,v_{x},v_{y},v_{z})}^{T}$. Using properties of normal distributions (see Ref. [eaton2007multivariate]) it can then be shown that the integrated coordinates conditioned on $x$ are also normally distributed, i.e., $\bm{\xi}_{I}|x\sim\mathcal{N}(\bm{\mu}_{\xi_{I}|x},\bm{\Sigma}_{\xi_{I}|x})$, with